Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 181     Entry time: 05 May 2023, 14:56
Author: Giuseppe Mercurio 
Group: SCS 
Subject: morning shift summary 

achievements

  • with help from LAS OCD, we got the PPL synchronized again with optical lock and close to the initial t0 on the sample
  • we found t0 on the sample again
  • collected data with magnetic field
  • initial setup of damnit/amore for semi-automatized data processing

problems

  • no issue besides the initial loss of PPL synchronization
ELOG V3.1.4-7c3fd00